OpenStudy (anonymous):

4x(x-1)-5x(x)=3 Solve for x

OpenStudy (anonymous):

\[4x(x-1)-5x(x)=3\] \[4x^2-4x-5x^2=3\] \[-x^2-4x-3=0\] \[x^2+4x+3=0\] \[(x+1)(x+3)=0\] \[x+1=0\] \[x=-1\] or \[x+3=0\] \[x=-3\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome